eviant:

PLS HELP!!! find the range of f(x)=-2x-5 for the domain (-2 -1 1 2)

dude:

To find the range, substitute each number in for x \(f(x)=-2(-2)-5\\ f(x)=-2(-1)-5\\ f(x)=-2(1)-5\\ f(x)=-2(2)-5\)

eviant:

thankx